I believe the answer is A but I would appreciate a second opinion.
Answer #1

Option 1.

Labor force = No of Unemployed + No of Employed

= 10+ No of Employed

= 10+40 = 50

As the unemployment rate is 20 percent, No of employed is 40 million

Unemployment rate = no of unemployed (No of employed + No of unemployed)

= 10/(1+40) = 10/50 =0.2 = 20 percent

Now with the working population of 70 million, and labor force equal to 50 million, people not in the labor force is 70-50=20 million